OpenStudy (anonymous):

What is the value of the expression 3–4? A. –81 B. –12 C. – 1/81 D. 1/81

OpenStudy (anonymous):

@radar @Nnesha

Nnesha (nnesha):

exponent rule \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

OpenStudy (danjs):

3^(-4)

OpenStudy (anonymous):

ok @Nnesha i have no time to explane plz give me answer plz i cant do this

OpenStudy (anonymous):

and yes

Nnesha (nnesha):

how do you know you `can't` do this

Nnesha (nnesha):

this is the easy topic in math

Nnesha (nnesha):

flip the fraction when you do that sign of exponent would change

Nnesha (nnesha):

here is an example \[\huge\rm 2^{-2} = \frac{ 1 }{ 2^2 }\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so \[3^{-4} =?\]

OpenStudy (anonymous):

1/18?

Nnesha (nnesha):

18 ?

OpenStudy (anonymous):

?\[\frac{ 1 }{ 18 }\]

Nnesha (nnesha):

not 18 3^4 is same as 3 times 3 times 3 times 3 which isn't equal to 18

Nnesha (nnesha):

\[\frac{ 1 }{ 3 \times 3 \times 3 \times 3}\]

OpenStudy (anonymous):

ok so -1/18?

OpenStudy (anonymous):

A. –81 B. –12 C. – 1/18 D. 1/18

Nnesha (nnesha):

3 times 3 times 3 times 3 = ??

OpenStudy (anonymous):

12

Nnesha (nnesha):

i'm serious.

Nnesha (nnesha):

http://prntscr.com/8biwdt these are answer choices

OpenStudy (anonymous):

yea

OpenStudy (mathstudent55):

To raise a number to a negative power, first, ignore the negative and just raise the number to the positive exponent. Then write 1 over the number. Examples: \(5^{-3} \) start with \(5^3 = 125\). Now write 1 over that. Answer \(\dfrac{ 1}{125} \) \(6^{-2}\) start with \(6^2 = 36\). Now write 1 over the number. Answer \(\dfrac{1}{36} \)

Nnesha (nnesha):

so 3 times 3 times 3 times 3 = what ?

OpenStudy (anonymous):

it is 81

OpenStudy (anonymous):

ok

Nnesha (nnesha):

yes right!!! so \[\frac{ 1 }{ 3 \times 3 \times 3 \times 3 }=\frac{ 1 }{ 81 }\] not 1/18

OpenStudy (anonymous):

ok so what next

OpenStudy (anonymous):

plz

Nnesha (nnesha):

that's it1

OpenStudy (anonymous):

ok and?

OpenStudy (anonymous):

what is answewr?

OpenStudy (anonymous):

er

Nnesha (nnesha):

seriously ?

OpenStudy (anonymous):

im just so furoustrated plz

Nnesha (nnesha):

reread my comments

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

lol, he has a slow streak

Nnesha (nnesha):

so what's the answer ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

but

OpenStudy (anonymous):

t

OpenStudy (anonymous):

it

OpenStudy (anonymous):

is

OpenStudy (anonymous):

not

OpenStudy (anonymous):

on

OpenStudy (anonymous):

my

OpenStudy (anonymous):

don't give him the answer if he is this lazy.

Nnesha (nnesha):

one is not answer pay attention reread my comments

Nnesha (nnesha):

like seriously ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz

Nnesha (nnesha):

I said read my comments again

OpenStudy (anonymous):

This guy is hopeless

OpenStudy (anonymous):

1/81

Nnesha (nnesha):

yes!

OpenStudy (anonymous):

shut up @MMAguy

Nnesha (nnesha):

thank you thank you!

OpenStudy (anonymous):

math is hard fo me

OpenStudy (anonymous):

do your own homework @shadybubba

OpenStudy (anonymous):

hes kinda right

Nnesha (nnesha):

when you don't put an effort then definitely it's hard

Nnesha (nnesha):

good luck!

OpenStudy (anonymous):

Nnesha such a hottie!

OpenStudy (anonymous):

waaaaaaaat?

OpenStudy (anonymous):

just help me

OpenStudy (anonymous):

no, you won't even help yourself.

OpenStudy (anonymous):

@Nnesha,@MMAguy likes you

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @shadybubba just help me \(\color{blue}{\text{End of Quote}}\) you got the answer for this question right!

OpenStudy (anonymous):

lawl @please_help_me

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

so answer is 1/18?

Nnesha (nnesha):

don't spam!

OpenStudy (anonymous):

81

Nnesha (nnesha):

1/81 not 18

OpenStudy (anonymous):

goodnes gracious look up and all see the answer!

OpenStudy (anonymous):

*u'll

OpenStudy (anonymous):

thx @Nnesha

Nnesha (nnesha):

you're welcome

OpenStudy (anonymous):

@MMAguy you were right

OpenStudy (anonymous):

about what? @please_help_me

OpenStudy (anonymous):

when you said: Nnesha such a hottie!

OpenStudy (anonymous):

jk lol

OpenStudy (anonymous):

lol